The length and the breadth of hall are in the ratio 5:4. If the breadth of the hall is 70m. Find its area?
Answer by laoman(51)

Given that the ratio of the length to the breadth is 5:4
The breadth is 70m
Let's assume that the length is 5x, so the breath is 4x
From these assumptions,
breath 4*x = 70
Therefore x = 70/4 = 17.5
Knowing x, we can get the length which is 5*x = 5*17.5 = 87.5
Now we know length to be 87.5m, breath to be 70m,
Area of rectangle = Length * breath
= 70m*87.5m = 6125 m^2
Dont forget the unit.
QED